### v1

@@ Abstract
Facial wrinkle detection plays a crucial role in cosmetic dermatology. Precise manual segmentation of facial wrinkles is challenging and time-consuming, with inherent subjectivity leading to inconsistent results among graders. To address this issue, we propose two solutions. First, we build and release the first public facial wrinkle dataset, ‘FFHQ-Wrinkle’, an extension of the NVIDIA FFHQ dataset. This dataset includes 1,000 images with human labels and 50,000 images with automatically generated weak labels. This dataset can foster the research community to develop advanced wrinkle detection algorithms. Second, we introduce a training strategy for U-Net-like encoder-decoder models to detect wrinkles across the face automatically. Our method employs a two-stage training strategy: texture map pretraining and finetuning on human-labeled data. Initially, we pretrain models on a large dataset with weak labels (N=50k) or masked texture maps generated through computer vision techniques, without human intervention. Subsequently, we finetune the models using human-labeled data (N=1k), which consists of manually labeled wrinkle masks. During finetuning, the network inputs a combination of RGB and masked texture maps, comprising four channels. We effectively combine labels from multiple annotators to minimize subjectivity in manual labeling. Our strategies demonstrate improved segmentation performance in facial wrinkle segmentation both quantitatively and visually compared to existing pretraining methods.
@@ 1 Introduction
Figure 1: Two stage training for facial wrinkle segmentation. (a) Weakly supervised pretraining stage: the model learns to extract masked texture maps from RGB face images. (b) Supervised finetuning stage: the model refines its ability to extract facial wrinkles from RGB masked face images and masked texture maps. The model parameters are initialized with the weights from the weakly supervised pretraining stage.
With the growing interest in dermatological diseases and skin aesthetics, predicting facial wrinkles is becoming increasingly significant. Facial wrinkles serve as critical indicators of aging [2, 19, 20], and are essential for evaluating skin conditions [29, 13], diagnosing dermatological disorders [30], and planning pre-treatment protocols for skin management [1, 27]. Nevertheless, the manual detection of facial wrinkles poses considerable challenges. Accurate detection and analysis of facial wrinkles necessitate a high level of expertise, typically available only through well-trained professionals such as dermatologists. This process is time-consuming and entails substantial costs due to the extensive time and effort required by the experts involved.
@@ 3.1 Dataset specifications
The first public facial wrinkle dataset, ‘FFHQ-Wrinkle’, comprises pairs of face images and their corresponding wrinkle masks. We focused on wrinkle labels while utilizing the existing high-resolution face image dataset FFHQ (Flickr-Faces-HQ) [12], which contains 70,000 high-resolution (1024x1024) face images captured under various angles and lighting conditions. The dataset we provide consists of one set of manually labeled wrinkle masks (N=1,000) and one set of "weak" wrinkle masks, or masked texture maps, generated without human labor (N=50,000). We selected 50,000 images from the FFHQ dataset, specifically image IDs 00000 to 49999. We used these 50,000 face images to create the weakly labeled wrinkles and randomly sampled 1,000 images from these to create the ground truth wrinkles. The methods for generating weakly labeled wrinkles and ground truth wrinkles are discussed in Section 4.2. As illustrated in Fig. 2, the dataset includes individuals of diverse ages, races, and backgrounds, featuring a variety of skin conditions such as freckles, acne, and pigmentation. This diversity makes the dataset particularly suitable for training models to handle the wide range of skin conditions encountered in clinical settings. The dataset is publicly available [^1] https://github.com/labhai/ffhq-wrinkle-dataset.
@@ 3.2 Ground truth wrinkle annotation
For ground truth wrinkles, we manually annotated the face images. The annotation process involved three annotators with extensive experience in image processing and analysis. Wrinkles can be categorized into two types—dynamic wrinkles and static wrinkles [31]. Dynamic wrinkles are formed by facial muscles and appear with expressions but disappear when the face is at rest. Static (permanent) wrinkles are visible even when the face is at rest and result from the repeated formation of dynamic wrinkles over time. Treatments for these categories also differ: dynamic wrinkles are often treated with muscle relaxants like Botox [28], while static wrinkles may require dermal fillers [8] or resurfacing treatments [7]. We annotated both types of wrinkles. Given the subjectivity inherent in wrinkle data, a consistent standard for wrinkle assessment was established prior to the commencement of labeling. The annotators conducted three synchronization sessions to minimize inter-rater variability. The annotation primarily targeted the forehead, crow’s feet, and nasolabial folds, encompassing the overall facial area. Due to the high resolution and diversity of the dataset—comprising various races, skin conditions, backgrounds, and angles—achieving consistent labeling results proved challenging, even with established standards for wrinkle assessment, as illustrated in Fig. 3. Consequently, as demonstrated in Table 1, the inter-rater agreement was low, underscoring the highly subjective nature of wrinkle assessments.
4 Method
4.1 Model architecture
@@ 4.2.1 Weakly supervised pretraining stage
where $G$ represents the Gaussian kernel, $\sigma$ denotes its standard deviation, $I_{G(\sigma)}$ is the Gaussian filtered image, and $(x,y)$ are the pixel coordinates in the image. Following the methodology in [14], we set the Gaussian kernel’s standard deviation to 5 and its size to 21x21 for texture map extraction. The extracted texture map contains detailed information about the contours, curves, and skin textures of the face image. However, as the texture map includes numerous false positives from the background, we employ a BiSeNet [35] architecture-based facial parsing deep learning model [^2] https://github.com/zllrunning/face-parsing.PyTorch to mask non-facial regions, resulting in the final masked texture map used as ground truth. We avoid converting the masked texture map into a binary mask by setting a threshold due to the variability in the size, shape, and depth of wrinkles, which makes determining an appropriate threshold challenging. Important wrinkle information can be lost during the binarization process, as illustrated in Fig. 5. Fig. 2-(b) shows the masked texture map used as the final ground truth in the weakly supervised pretraining stage.
@@
Figure 5: Loss of wrinkle information due to binarization. Comparison of the masked texture map with the binary masked texture map obtained through Otsu’s binarization [22] method. (a) Face image. (b) Masked texture map. (c) Binary masked texture map.
@@ 5.4 Ablation study
Swin UNETR with pretraining None 0.4220 0.5858 0.9949 0.2545 0.3944 0.9932 25.153M
ImageNet-1K [24] (Swin-T [16]) 0.4385 0.6028 0.9952 0.2877 0.4351 0.9939 100.557M
Masked image prediction [32] 0.4450 0.6079 0.9954 0.2963 0.4452 0.9937 25.153M
@@
U-Net with pretraining None 0.4638 0.6278 0.9955 0.3021 0.4551 0.9918 17.263M
@@ 6 Discussion
Our approach achieves state-of-the-art performance when compared to two publicly released models specifically designed for wrinkle segmentation, in addition to outperforming ImageNet pretrained models and self-supervised learning methods. We demonstrate that our two-stage training strategy significantly enhances wrinkle segmentation efficiency. Furthermore, our approach shows the potential to achieve high performance with limited data, which could enhance scalability and flexibility in clinical settings. By using a large amount of weakly labeled data obtained automatically through filters for weakly supervised training and then finetuning with a small amount of reliable manually labeled data, we significantly reduce the time and cost required for manual labeling while improving the segmentation performance of facial wrinkles. To minimize subjectivity in the manual labeling process, we effectively combine data labeled by multiple annotators, resulting in more reliable training data. Additionally, to enhance the reproducibility of our research and reduce the manual labeling costs for subsequent studies, we release the dataset publicly available, which can also serve as a benchmark dataset for future research. The performance improvement of facial wrinkle segmentation through transfer learning has not been conducted in previous research, indicating that our approach can be efficiently integrated into various tasks related to facial wrinkle detection and segmentation tasks.
@@
However, our approach has limitations. As shown in Fig. 8, objects similar to wrinkles, such as hair or fingers covering the face, are mistakenly recognized as wrinkles in the images. This results in false positives (FP) during the wrinkle segmentation process. To address this issue, future research focuses on developing techniques that can accurately segment facial regions and precisely distinguish between wrinkle and non-wrinkle areas to reduce FP. Additionally, despite majority voting, the subjectivity in wrinkle annotation remains a challenge. In future research, we will collaborate with dermatologists for wrinkle annotation and explore techniques such as label smoothing to improve the reliability and trustworthiness of ground truth wrinkles.
@@ 7 Conclusion
We propose a two-stage learning strategy for facial wrinkle segmentation using deep learning. Our approach leverages the knowledge gained from extracting facial features for transfer learning in facial wrinkle segmentation. This allows us to achieve high segmentation performance even with limited manually labeled data, thereby establishing a state-of-the-art training strategy. This significantly reduces the time and cost required for manual labeling of wrinkles, thus offering potential benefits in cosmetic dermatology. Furthermore, to support ongoing research and enhance reproducibility, we have made our facial wrinkle dataset publicly available.
@@ 7.0.1 Acknowledgements
The authors appreciate Dr. Ik Jun Moon, a dermatologist at Asan Medical Center, for sharing invaluable insights and feedback from a dermatological perspective. This work was supported by the National Supercomputing Center with supercomputing resources including technical support (KSC-2024-CRE-0021), Hankuk University of Foreign Studies Research Fund of 2024, the National Research Foundation of Korea (NRF) grant funded by the Ministry of Science and ICT (MSIT) (RS-2024-00338048), the artificial intelligence semiconductor support program to nurture the best talents (IITP(2024)-RS-2023-00253914) grant funded by the Korea government, and the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024(RS-2024-00332210).

### v4

@@ Abstract
Facial wrinkle detection plays a crucial role in cosmetic dermatology. Precise manual segmentation of facial wrinkles is challenging and time-consuming, with inherent subjectivity leading to inconsistent results among graders. To address this issue, we propose two solutions. First, we build and release the first public facial wrinkle dataset, ‘FFHQ-Wrinkle’, an extension of the NVIDIA FFHQ dataset. It includes 1,000 images with human labels and 50,000 images with automatically generated weak labels. This dataset could serve as a foundation for the research community to develop advanced wrinkle detection algorithms. Second, we introduce a simple training strategy utilizing texture maps, applicable to various segmentation models, to detect wrinkles across the face. Our two-stage training strategy first pretrain models on a large dataset with weak labels (N=50k), or masked texture maps generated through computer vision techniques, without human intervention. We then finetune the models using human-labeled data (N=1k), which consists of manually labeled wrinkle masks. The network takes as input a combination of RGB and masked texture map of the image, comprising four channels, in finetuning. We effectively combine labels from multiple annotators to minimize subjectivity in manual labeling. Our strategies demonstrate improved segmentation performance in facial wrinkle segmentation both quantitatively and visually compared to existing pretraining methods. The dataset is available at https://github.com/labhai/ffhq-wrinkle-dataset.
@@ 1 Introduction
Figure 1: Two-stage training for facial wrinkle segmentation. (a) Weakly supervised pretraining stage: the model learns to extract masked texture maps from RGB face images. (b) Supervised finetuning stage: the model refines its ability to extract facial wrinkles from RGB-masked face images and masked texture maps. The model parameters are initialized with the weights from the weakly supervised pretraining stage.
With the growing interest in dermatological diseases and skin aesthetics, predicting facial wrinkles is becoming increasingly significant. Facial wrinkles serve as critical indicators of aging [2, 19, 20], and are essential for evaluating skin conditions [29, 13], diagnosing dermatological disorders [30], and planning pre-treatment protocols for skin management [1, 27]. Nevertheless, the manual detection of facial wrinkles poses considerable challenges. Accurate detection and analysis of facial wrinkles necessitate a high level of expertise, typically available only through well-trained professionals such as dermatologists. This process is time-consuming and entails substantial costs due to the extensive time and effort required by the experts.
@@ 3.1 Dataset specifications
The first public facial wrinkle dataset, ‘FFHQ-Wrinkle’, comprises pairs of face images and their corresponding wrinkle masks. We focused on wrinkle labels while utilizing the existing face image dataset FFHQ (Flickr-Faces-HQ) [12], which contains 70,000 high-resolution (1024x1024) face images captured under various angles and lighting conditions. The dataset we provide consists of one set of manually labeled wrinkle masks (N=1,000) and one set of ‘weak’ wrinkle masks, or masked texture maps, generated without human labor (N=50,000). We selected 50,000 images from the FFHQ dataset, specifically image IDs 00000 to 49999. We used these 50,000 face images to create the weakly labeled wrinkles and randomly sampled 1,000 images from these to create the ground truth wrinkles. The methods for generating weakly labeled wrinkles and ground truth wrinkles are discussed in Section 4.2. Table 1 summarizes estimated demographic information of the dataset–i.e. age, race, and sex. The age and sex data were sourced from the FFHQ-Aging [22] dataset, where at least three annotators labeled each image. The race/ethnicity attribute was obtained through facial attribute analysis using the DeepFace [^1] https://github.com/serengil/deepface framework. Hence, the demographic information may include errors. As illustrated in Fig. 2, the dataset consists of individuals of varying ages, sex, and race/ethnicity, featuring a range of skin conditions such as freckles, acne, and pigmentation. This diversity makes the dataset particularly suitable for training models to handle the wide array of skin conditions encountered in clinical settings. The dataset is publicly available at https://github.com/labhai/ffhq-wrinkle-dataset.
Table 1: Demographic attributes of the dataset. The ‘Human-labeled’ data represents the 1,000 face images manually labeled by human annotators and the ‘Weakly-labeled’ data refers to the 50,000 images labeled without human intervention.
Dataset Human-labeled Weakly-labeled
Sample size 1000 50000
Age 0-9 / 10-19 / 20-29 / 66 / 68 / 233 /   7030 / 4448 / 13804 /
30-39 / 40-49 / 50-69 / 70+   246 / 186 / 161 / 40   10960 / 6931 / 5550 / 1277
Sex Male / Female 471 / 529 26929 / 23071
Race/ Ethnicity   White / Asian / Latino Hispanic / 587 / 210 / 67 / 29728 / 11121 / 3895 /
Black / Middle Eastern / Indian 81 / 37 / 18 / 2383 / 2053 / 820
@@ 3.2 Ground truth wrinkle annotation
For ground truth wrinkles, we manually annotated the face images. The annotation process involved three annotators with extensive experience in image processing and analysis. Wrinkles can be categorized into two types—dynamic wrinkles and static wrinkles [31]. Dynamic wrinkles are formed by facial muscles and appear with expressions but disappear when the face is at rest. Static (permanent) wrinkles are visible even when the face is at rest and result from the repeated formation of dynamic wrinkles over time. We annotated both types of wrinkles without distinguishing between them. Given the subjectivity inherent in wrinkle data, a consistent standard for wrinkle assessment was established prior to the commencement of labeling. The annotators conducted three synchronization sessions to minimize inter-rater variability. The annotation primarily targeted the forehead, crow’s feet, and nasolabial folds, encompassing the overall facial area. Due to the high resolution and diversity of the dataset—comprising various races, skin conditions, backgrounds, and angles—achieving consistent labeling results proved challenging, even with established standards for wrinkle assessment, as illustrated in Fig. 3. Consequently, as demonstrated in Table 2, the inter-rater agreement was low, underscoring the highly subjective nature of wrinkle assessments.
@@ 4.2.1 Weakly supervised pretraining stage
where $G$ represents the Gaussian kernel, $\sigma$ denotes its standard deviation, $I_{G(\sigma)}$ is the Gaussian filtered image, and $(x,y)$ are the pixel coordinates in the image. Following the methodology in [14], we set the Gaussian kernel’s standard deviation to 5 and its size to 21x21 for texture map extraction. The extracted texture map contains detailed information about the contours, curves, and skin textures of the face image. However, as the texture map includes numerous false positives from the background, we employ a BiSeNet [35] architecture-based facial parsing deep learning model [^2] https://github.com/zllrunning/face-parsing.PyTorch to mask non-facial regions, resulting in the final masked texture map used as ground truth. We avoid converting the masked texture map into a binary mask due to the variability in the size, shape, and depth of wrinkles, which makes determining an appropriate threshold challenging. Fig. 2-(b) shows the masked texture map used as the final ground truth in the weakly supervised pretraining stage.
@@ 5.4 Ablation study
Swin UNETR with pretraining No pretraining 0.4220 0.5858 0.9949 0.2545 0.3944 0.9932 25.153M
ImageNet-1K [24] (Swin-T [16]) 0.4385 0.6028 0.9952 0.2877 0.4351 0.9939 100.56M
Masked image modeling [32] 0.4450 0.6079 0.9954 0.2963 0.4452 0.9937 25.153M
@@
U-Net with pretraining No pretraining 0.4638 0.6278 0.9955 0.3021 0.4551 0.9918 17.263M
@@ 6 Discussion
Our approach achieves state-of-the-art performance when compared to two publicly released models specifically designed for wrinkle segmentation, in addition to outperforming ImageNet pretrained models and self-supervised learning methods. We demonstrate that our two-stage training strategy significantly enhances wrinkle segmentation efficiency. Furthermore, our approach shows the potential to achieve high performance with limited data, which could enhance scalability and flexibility in clinical settings. By using a large amount of weakly labeled data obtained automatically through filters for weakly supervised training and then finetuning with a small amount of reliable manually labeled data, we significantly reduce the time and cost required for manual labeling while improving the segmentation performance of facial wrinkles. To minimize subjectivity in the manual labeling process, we effectively combine data labeled by multiple annotators, resulting in more reliable training data. Additionally, to enhance the reproducibility of our research and reduce the manual labeling costs for subsequent studies, we release the dataset publicly available, which can also serve as a benchmark dataset for future research. The performance improvement of facial wrinkle segmentation through transfer learning has not been conducted in previous research, indicating that our approach can be efficiently integrated into various tasks related to facial wrinkle detection and segmentation tasks. Additionally, since this research falls under the broader category of thin object detection tasks, it is expected to be widely applicable to studies requiring segmentation of thin objects (e.g., fundus imaging, vascular imaging).
@@
However, our approach has limitations. As shown in Fig. 7, objects similar to wrinkles, such as hair or fingers covering the face, are mistakenly recognized as wrinkles in the images. This results in false positives during the wrinkle segmentation process. To address this issue, upcoming studies will focus on developing techniques that can accurately segment facial regions and precisely distinguish between wrinkle and non-wrinkle areas to reduce false positives. Also, there may be benefits to including the type of wrinkle (e.g., static vs. dynamic wrinkle) to each wrinkle in the facial image because treatment strategies often differ by the type in clinics [28, 8, 7]. Despite majority voting, the subjectivity in wrinkle annotation remains a challenge. Moving forward, we plan to collaborate with dermatologists for wrinkle annotation and explore techniques such as soft labeling to improve the reliability and trustworthiness of ground truth wrinkles.
@@ 7 Conclusion
We propose a two-stage learning strategy for facial wrinkle segmentation that leverages transfer learning from facial texture feature extraction. Specifically, the model is pretrained using automatically generated weak wrinkle labels (masked texture maps) to learn general facial features such as contours and skin texture. The model is then finetuned with a smaller set of manually labeled wrinkle data to enhance segmentation performance. This method demonstrates both qualitatively and quantitatively superior results, achieving state-of-the-art performance. Consequently, it significantly reduces the time and cost of manual wrinkle labeling, offering potential benefits in cosmetic dermatology. Additionally, the pretraining method’s architecture-independent nature suggests its broad applicability to various segmentation models, making it valuable not only in facial wrinkle segmentation but also in other areas requiring the segmentation of thin objects where manual labeling is costly. To support ongoing research and reproducibility, we have made the FFHQ-Wrinkle dataset—the first publicly available dataset of its kind—accessible to the research community. This dataset comprises 1,000 manually labeled wrinkle images and 50,000 weakly labeled images. By sharing this dataset, we aim to facilitate the development of more advanced wrinkle detection models and promote further advancements in this field.
@@ 7.0.1 Acknowledgements
The authors appreciate Dr. Ik Jun Moon, a dermatologist at Asan Medical Center, for sharing invaluable insights and feedback from a dermatological perspective. This work was supported by the National Research Foundation of Korea (NRF) grants funded by the Ministry of Science and ICT (MSIT) (RS-2024-00455720 & RS-2024-00338048), the National Institute of Health(NIH) research project (2024ER040700), the National Supercomputing Center with supercomputing resources including technical support (KSC-2024-CRE-0021), Hankuk University of Foreign Studies Research Fund of 2024, the artificial intelligence semiconductor support program to nurture the best talents (IITP(2024)-RS-2023-00253914) grant funded by the Korea government, and the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism in 2024(RS-2024-00332210).